# Optimizing Robustness and Accuracy in Mixture of Experts: A Dual-Model Approach

### v2

@@ Abstract
Mixture of Experts (MoE) have shown remarkable success in leveraging specialized expert networks for complex machine learning tasks. However, their susceptibility to adversarial attacks presents a critical challenge for deployment in robust applications. This paper addresses the critical question of how to incorporate robustness into MoEs while maintaining high natural accuracy. We begin by analyzing the vulnerability of MoE components, finding that expert networks are notably more susceptible to adversarial attacks than the router. Based on this insight, we propose a targeted robust training technique that integrates a novel loss function to enhance the adversarial robustness of MoE, requiring only the robustification of one additional expert without compromising training or inference efficiency. Building on this, we introduce a dual-model strategy that linearly combines a standard MoE model with our robustified MoE model using a smoothing parameter. This approach allows for flexible control over the robustness-accuracy trade-off. We further provide theoretical foundations by deriving certified robustness bounds for both the single MoE and the dual-model. To push the boundaries of robustness and accuracy, we propose a novel joint training strategy JTDMoE for the dual-model. This joint training enhances both robustness and accuracy beyond what is achievable with separate models. Experimental results on CIFAR-10 and TinyImageNet datasets using ResNet18 and Vision Transformer (ViT) architectures demonstrate the effectiveness of our proposed methods.
@@ 3 Preliminaries
Adversarial Attacks on MoE Models. An adversarial attack on the entire MoE model seeks to find a perturbation ${\bm{\delta}}$ such that the model’s output on the perturbed input ${\mathbf{x}}+{\bm{\delta}}$ differs significantly from the output on the original input ${\mathbf{x}}$ or its ground truth label $y$: $F({\mathbf{x}}+{\bm{\delta}})=\sum_{i=1}^{E}a_{i}({\mathbf{x}}+{\bm{\delta}})f_{i}({\mathbf{x}}+{\bm{\delta}})$, where the ${\bm{\delta}}$ is usually generated by maximizing the loss function $\arg\max_{\|{\bm{\delta}}\|_{p}\leq\epsilon}{\mathcal{L}}(F({\mathbf{x}}+{\bm{\delta}}),y)$ under the $\ell_{p}$ norm and the attack budget $\epsilon$. We also consider loss terms that do not rely on $y$. Notice that the attack is on the whole MoE. We implement adversarial attacks focusing on specific components of the MoE model. An attack targeting the router aims to alter the routing weights without affecting the experts’ outputs. The perturbation ${\bm{\delta}}$ is crafted by changing the router’s decisions while the experts receiving the original input: $F({\mathbf{x}}+{\bm{\delta}})=\sum_{i=1}^{E}a_{i}({\mathbf{x}}+{\bm{\delta}})f_{i}({\mathbf{x}})$. An attack targeting the experts aims to alter the experts’ outputs without changing the routing weights. ${\bm{\delta}}$ is designed such that: $F({\mathbf{x}}+{\bm{\delta}})=\sum_{i=1}^{E}a_{i}({\mathbf{x}})f_{i}({\mathbf{x}}+{\bm{\delta}})$. We remark that once the perturbation ${\bm{\delta}}$ is generated, it will be implemented on both the router and experts despite of its target.
@@ 4.2 Incorporating Robustness into MoE Architecture
where $f_{2}({\mathbf{x}})$ denotes the output of the expert with the second-largest router weight for the adversarial example. The term $\ell_{KL}(f_{2}({\mathbf{x}}+{\bm{\delta}}),f_{2}({\mathbf{x}}))$ represents the KL divergence between the output of this expert on the adversarial example and its output on the clean input. RT-ER is compatible with various routing strategies. For instance, when the router employs a top- n strategy, RT-ER selects an additional expert—one not initially chosen by the router—based on the router weights, leveraging the router’s output. We use KL divergence instead of cross-entropy because cross-entropy requires a well-defined ground truth for the target distribution, which is not applicable in this context. The hyperparameter $\beta$ controls the trade-off between MoE-wide robustness and expert-specific robustness. For clarity, we refer to the expert with the second-largest router weight as the second-top expert in the following discussion.
@@ 5.1 Dual-Model with Post-Training MoE
Based on Theorem 5.5, there are three ways to further enhance the robustness of the dual-model: increasing the value of $\alpha$, enlarging the margin, and decreasing the Lipschitz constant of both the experts and the router. To ensure that the dual-model retains its robustness, $\alpha$ must remain within the range of $[0.5,1]$. Furthermore, a higher value of $\alpha$, which implies a larger contribution of the robust MoE to the final prediction, would result in a decrease in SA, which contradicts the motivation for using the dual-model architecture. As for the Lipschitz constants, they have already been optimized during RT-ER. In summary, the most efficient way to enhance the dual-model’s performance is by enlarging its margin. Additionally, a larger margin also leads to better accuracy on clean data.
@@ 6.1 Experiment Setup
Our experiments use the MoE architecture on CIFAR-10 (Krizhevsky et al., 2009) and TinyImageNet (Deng et al., 2009), with a fully connected top-1 router and 4 experts ($E=4$). We use ResNet18 (He et al., 2016) for CIFAR-10 and pre-trained ViT-small (Liu et al., 2021) for TinyImageNet. Instead of training the dual-model from scratch, we apply the JTDMoE algorithm to pre-trained models for efficiency, using ST MoE as the standard and RT-ER MoE as the robust MoE, with $\alpha$ ranging from 0.5 to 1.
@@ 7 Conclusion
Our study presents a framework for enhancing adversarial robustness and accuracy in MoE architectures, focusing on vulnerable expert networks and strategies that balance both, supported by theoretical analysis and empirical validation.
8 Impact Statement
@@ Appendix A Supplementary Material
In this supplementary material, we provide the proofs of Theorems 5.4 and 5.5 in Section A.1. Additional experimental results are presented as follows: single MoE experiments in Section A.2, dual-model experiments based on pre-trained MoE in Section A.3, and JTDMoE experiments in Section A.4.
@@ A.1 Proof of Key Theorems
A.2 Additional Experiments on Robust MoE
Performance of RT-ER on TinyImageNet Dataset
For For the TinyImageNet dataset (Deng et al., 2009), we adversarially trained an MoE model using ViT-small as the experts. The experimental setup follows the same settings outlined in Section 6. The RA and SA performance curves are shown in Figure 4.
@@ A.2 Additional Experiments on Robust MoE
Comparison with Trades and AdvMoE
@@
The Impact of Number of Experts $E$
@@ E SA(%) RA(%) RA-E(%) RA-R(%)
The Impact of Different $\beta$
@@
A.3 Experiments with Pre-trained Dual-Model
@@
A.4 JTDMoE Experimental Results
Effectiveness of JTDMoE Under Different $\alpha$
@@ A.4 JTDMoE Experimental Results
CIFAR-10
Class Airplane Automobile Bird Cat Deer
Improvement 5.11% 3.49% 7.5% 5.55% 10.27%
Class Dog Frog Horse Ship Truck
Improvement 14.42% 6.46% 4.03% 1.25% 4.58%
Margin Comparison

### v3

@@ Abstract
Mixture of Experts (MoE) have shown remarkable success in leveraging specialized expert networks for complex machine learning tasks. However, their susceptibility to adversarial attacks presents a critical challenge for deployment in robust applications. This paper addresses the critical question of how to incorporate robustness into MoEs while maintaining high natural accuracy. We begin by analyzing the vulnerability of MoE components, finding that expert networks are notably more susceptible to adversarial attacks than the router. Based on this insight, we propose a targeted robust training technique that integrates a novel loss function to enhance the adversarial robustness of MoE, requiring only the robustification of one additional expert without compromising training or inference efficiency. Building on this, we introduce a dual-model strategy that linearly combines a standard MoE model with our robustified MoE model using a smoothing parameter. This approach allows for flexible control over the robustness-accuracy trade-off. We further provide theoretical foundations by deriving certified robustness bounds for both the single MoE and the dual-model. To push the boundaries of robustness and accuracy, we propose a novel joint training strategy JTDMoE for the dual-model. This joint training enhances both robustness and accuracy beyond what is achievable with separate models. Experimental results on CIFAR-10 and TinyImageNet datasets using ResNet18 and Vision Transformer (ViT) architectures demonstrate the effectiveness of our proposed methods. The code is publicly available at https://github.com/TIML-Group/Robust-MoE-Dual-Model.
@@ 3 Preliminaries
Adversarial Attacks on MoE Models. An adversarial attack on the entire MoE model seeks to find a perturbation ${\bm{\delta}}$ such that the model’s output on the perturbed input ${\mathbf{x}}+{\bm{\delta}}$ differs significantly from the output on the original input ${\mathbf{x}}$ or its ground truth label $y$: $F({\mathbf{x}}+{\bm{\delta}})=\sum_{i=1}^{E}a_{i}({\mathbf{x}}+{\bm{\delta}})f_{i}({\mathbf{x}}+{\bm{\delta}})$, where the ${\bm{\delta}}$ is usually generated by maximizing the loss function $\arg\max_{\|{\bm{\delta}}\|_{p}\leq\epsilon}{\mathcal{L}}(F({\mathbf{x}}+{\bm{\delta}}),y)$ under the $\ell_{p}$ norm and the attack budget $\epsilon$. We also consider loss terms that do not rely on $y$. Notice that the attack is on the whole MoE. We implement adversarial attacks focusing on specific components of the MoE model. An attack targeting the router aims to alter the routing weights without affecting the experts’ outputs. The perturbation ${\bm{\delta}}$ is crafted by changing the router’s decisions while the experts receiving the original input: $F({\mathbf{x}}+{\bm{\delta}})=\sum_{i=1}^{E}a_{i}({\mathbf{x}}+{\bm{\delta}})f_{i}({\mathbf{x}})$. An attack targeting the experts aims to alter the experts’ outputs without changing the routing weights. ${\bm{\delta}}$ is designed such that: $F({\mathbf{x}}+{\bm{\delta}})=\sum_{i=1}^{E}a_{i}({\mathbf{x}})f_{i}({\mathbf{x}}+{\bm{\delta}})$. We remark that once the perturbation ${\bm{\delta}}$ is generated, it will be implemented on both the router and experts despite of its target. When attacking the expert networks, the perturbations are generated under the assumption that the routing scores remain unchanged. The generated perturbations can still alter the final predictions of MoE models, as they are largely independent of the router’s behavior. Empirically, our experiments show that in 98% of cases, the perturbed inputs are still routed to the same expert(s), indicating that the expert-targeted attacks remain effective despite this assumption.
@@ 4.2 Incorporating Robustness into MoE Architecture
where $f_{2}({\mathbf{x}})$ denotes the output of the expert with the second-largest router weight for the adversarial example. The term $\ell_{KL}(f_{2}({\mathbf{x}}+{\bm{\delta}}),f_{2}({\mathbf{x}}))$ represents the KL divergence between the output of this expert on the adversarial example and its output on the clean input. RT-ER is compatible with various routing strategies. For instance, when the router employs a top- n strategy, RT-ER selects an additional expert—one not initially chosen by the router—based on the router weights, leveraging the router’s output. We adopt KL divergence instead of cross-entropy for two primary reasons. First, cross-entropy relies on a well-defined ground-truth target distribution, which is unavailable in our setting. Second, employing cross-entropy in the second term would compel the predictions to closely match the ground truth, potentially resulting in overfitting. In contrast, KL divergence enables a softer alignment between experts, thereby reducing the risk of overfitting and offering greater flexibility in modeling uncertainty. The hyperparameter $\beta$ controls the trade-off between MoE-wide robustness and expert-specific robustness. For clarity, we refer to the expert with the second-largest router weight as the second-top expert in the following discussion.
@@ 5.1 Dual-Model with Post-Training MoE
Our theoretical results are all based on Assumption 5.3, which constrains the applicability of our method. Therefore, it is important to clarify the scope under which this assumption holds. We identify three representative scenarios where Assumption 5.3 naturally applies. First, in sparse MoE after robust training, robust optimization techniques typically encourage large margins in expert scores. This leads to a clear separation between the top-n experts and the rest, making it difficult for small adversarial perturbations to change the top-n set. As a result, the router becomes locally stable and selects the same expert(s) for both clean and perturbed inputs. Second, in dense MoE (Zhang et al., 2024), all experts are activated and the routing function is continuous, which trivially satisfies the assumption. Third, soft MoE (Puigcerver et al., 2023) also employs a continuous routing function with soft assignments, ensuring that small input changes lead to smooth changes in routing outputs. In all these cases, the routing behavior is stable under small perturbations, thus meeting the requirement of Assumption 5.3.
@@
Based on Theorem 5.5, there are three ways to further enhance the robustness of the dual-model: increasing the value of $\alpha$, enlarging the margin, and decreasing the Lipschitz constant of both the experts and the router. The maximum value of $F_{R}^{(y)}(\mathbf{x})-F_{R}^{(k)}(\mathbf{x})$ is 1. If $\alpha$ falls below 0.5, the numerator in the certified robustness bound becomes negative, rendering the robustness radius undefined. Therefore, to ensure that the dual-model maintains a valid certified robustness guarantee, $\alpha$ must lie within the range $[0.5,1]$. Furthermore, a higher value of $\alpha$, which implies a larger contribution of the robust MoE to the final prediction, would result in a decrease in SA, which contradicts the motivation for using the dual-model architecture. As for the Lipschitz constants, they have already been optimized during RT-ER. In summary, the most efficient way to enhance the dual-model’s performance is by enlarging its margin. Additionally, a larger margin also leads to better accuracy on clean data.
@@ 6.1 Experiment Setup
Our experiments use the MoE architecture on CIFAR-10 (Krizhevsky et al., 2009) and TinyImageNet (Deng et al., 2009), with a fully connected top-1 router and 4 experts ($E=4$). We use ResNet18 (He et al., 2016) for CIFAR-10 and pre-trained ViT-small (Liu et al., 2021) for TinyImageNet. Instead of training the dual-model from scratch, we apply the JTDMoE algorithm to pre-trained models for efficiency, using ST MoE as the standard and RT-ER MoE as the robust MoE, with $\alpha$ ranging from 0.5 to 1. More details can be found in Appendix A.2.
@@ 6.4 Evaluation of JTDMoE
Margin Comparison.
CIFAR-10
Class Airplane Automobile Bird Cat Deer
Improvement 5.11% 3.49% 7.5% 5.55% 10.27%
Class Dog Frog Horse Ship Truck
Improvement 14.42% 6.46% 4.03% 1.25% 4.58%
@@ 7 Conclusion
This work presents a comprehensive framework for enhancing adversarial robustness in MoE models. Our approach is motivated by the key observation that expert networks in MoEs are significantly more susceptible to adversarial perturbations than the router—a structural vulnerability specific to the MoE architecture. To address this, we propose a targeted robustification strategy with accompanying theoretical robustness bounds, which adversarially trains an additional expert not selected by the router. Building upon the theoretical robustness–accuracy trade-off, we further introduce a dual-model framework (JTDMoE) that integrates a standard MoE and a robust MoE via a bi-level joint training scheme, achieving strong adversarial robustness without compromising standard accuracy. Extensive experiments across diverse models and datasets validate the effectiveness and scalability of our proposed methods.
Impact Statement
@@
Acknowledgement
This work is supported by the NSF under Grants 2246157 and 2319243. We are thankful for the computational resources made available through NSF ACCESS and Argonne Leadership Computing Facility.
@@ Appendix A Supplementary Material
In this supplementary material, we provide the proofs of Theorems 5.4 and 5.5 in Section A.1. Additional experimental results are organized as follows: single MoE experiments are reported in Section A.3, dual-model experiments based on pre-trained MoEs are presented in Section A.4, and JTDMoE experiments are detailed in Section A.5. In addition, Section A.6 presents RT-ER experiments conducted using a large model (ViT) on the large-scale ImageNet dataset.
@@ A.1 Proof of Key Theorems
A.2 Model Architecture Details
Following several recent works (Videau et al., 2024; Chen et al., 2025; He et al., 2024), the default setting is a single MoE layer as a replacement for the classification layer. We further conduct experiments using the architecture proposed by Riquelme et al. (2021), which inserts MoE layers in place of the dense MLPs within transformer blocks. After standard training, the MoE model achieves 90.35% SA, 38.02% RA, 32.16% RA-E, and 64.97% RA-R. These results further support our observation that expert networks are generally more vulnerable to adversarial perturbations than the router. This vulnerability likely stems from the fact that expert networks are deeper and more complex than the router, making them inherently more susceptible to such attacks. Regardless of the specific MoE architecture used, this trend remains consistent: the expert networks exhibit greater vulnerability due to their complexity.
A.3 Additional Experiments on Robust MoE
Performance of RT-ER on TinyImageNet Dataset.
For the TinyImageNet dataset (Deng et al., 2009), we adversarially trained an MoE model using ViT-small as the experts. The experimental setup follows the same settings outlined in Section 6. The RA and SA performance curves are shown in Figure 4.
@@ A.3 Additional Experiments on Robust MoE
(a) SA
(b) RA
Comparison with Trades and AdvMoE.
@@
The Impact of Number of Experts $E$.
@@ E SA(%) RA(%) RA-E(%) RA-R(%)
The Impact of Different $\beta$.
@@
A.4 Experiments with Pre-trained Dual-Model
@@
A.5 JTDMoE Experimental Results
Effectiveness of JTDMoE Under Different $\alpha$.
@@ A.5 JTDMoE Experimental Results
A.6 ViT Experimental results on ImageNet
Performance of RT-ER on the ImageNet Dataset.
To evaluate the scalability of RT-ER to large-scale models and datasets, we conduct experiments using ViT on the ImageNet dataset (Deng et al., 2009). The experimental setup follows the protocol described in Section 6, and the results are summarized in Table 10. RT-ER consistently outperforms both AT and TRADES across all evaluation metrics—SA (+8%), RA (+12%), RA-E (+1%), and RA-R (+0.6%)—demonstrating its effectiveness and efficiency in improving the robustness of MoE models at scale.
Table 10: Robustness evaluation of AT, TRADES, and RT-ER. RT-ER achieves a substantial improvement of approximately 12% in RA and 8% in SA compared to conventional adversarial training. The evaluation is conducted on the ImageNet dataset using a MoE model with ViT-based experts.
ImageNet
Method SA(%) RA(%) RA-E(%) RA-R(%)
AT 60.32 44.64 43.06 70.24
TRADES 61.94 45.54 43.75 70.37
RT-ER 68.38 56.16 44.99 70.82